Periodic Inventory Using FIFO, LIFO, and Weighted Average Cost Method. The units of an item available for sale during the year were as follows: Jan. 1 Inventory 11 units at $45 Aug. 13 Purchase 18 units at $46 Nov. 30 Purchase Available for sale There are 16 units of the item in the physical inventory at December 31. The periodic inventory system is used. Determine the inventory cost using (a) the first-in, first-out (FIFO) method; (b) the last-in, first-out (LIFO) method; and (c) the weighted average cost method (round per-unit cost to two decimal places and your final answer to the nearest whole dollar). a. First-in, first-out (FIFO) $495 828 480 $1,803 10 units at $48 39 units b. Last-in, first-out (LIFO) c. Weighted average cost
Answer #1

Calculate ending inventory :

a) First in First out = (10*48+6*46) = 756

b) Last in first out = (11*45+5*46) = 725

c) Weighted average method = (1803/39*16) = 740